OpenStudy (laddiusmaximus):

solve using squeeze theorem limx-->0 x sin 1/x^2

OpenStudy (turingtest):

just as zarkon did with the previous problem\[-x\le x\sin(\frac1{x^2})\le x\]

OpenStudy (laddiusmaximus):

thats it?